OpenStudy (anonymous):

A pyramid has a square base with sides of length s. The height of the pyramid is equal to 1/2 of the length of a side on the base. Which formula represents the volume of the pyramid?

OpenStudy (anonymous):

(A) V=1/12s^3 (B) V=1/6s^3 (C) V=1/3s^3 (D) V=3s^3 (E) V=6s^3

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

@PeterPan could you be of help on this problem.

OpenStudy (anonymous):

I remember the Ancient Egyptians doing this sort of thing :D The volume of a pyramid is equal to one third of the product of its length, width, and height. What say you? ^.^

OpenStudy (anonymous):

So it is C?

OpenStudy (anonymous):

how did you get C?

OpenStudy (anonymous):

It has 1/3 in it.

OpenStudy (anonymous):

haha they all have 1/3 in them, hidden somewhere XD Look, what are the length, width, and height of this pyramid? :)

OpenStudy (anonymous):

width is = to s. the height is 1/2 of s. Not sure about the base.

OpenStudy (anonymous):

i mean length

OpenStudy (anonymous):

Not sure about length, but you're sure about width? Silly Mr. Man, it's a square base ^.^

OpenStudy (anonymous):

It is isn't it haha.

OpenStudy (anonymous):

It makes things so much easier when they give a picture. Im a better when i can see it with my eyes.

OpenStudy (anonymous):

Okay, length is s, width is s, height is half of s. Multiply these three, you get...what?

OpenStudy (anonymous):

Is sxs 2s or s^2

OpenStudy (anonymous):

Oh no :O

OpenStudy (anonymous):

You're so lost.... lost boy (man?) LOL sxs = s^2 s+s = 2s know the difference, mister :P

OpenStudy (anonymous):

I am lost (: thanks for puttin up with this.

OpenStudy (anonymous):

i've put up with worse..... ;) now multiply s, s, and \(\large \frac12s\)

OpenStudy (anonymous):

kk sxs=s^2 then i multiply that by .5, but i get 1. so would it just be (s)?

OpenStudy (anonymous):

maybe this class is a little too hard? :3

OpenStudy (anonymous):

well, for now, anyway :D There's just so many things wrong with your logic, that I'm trying to find where to begin... now wait a sec...

OpenStudy (anonymous):

Regardless if you guide me to my answer or not. You'll be receiving a nice little testimonial.

OpenStudy (anonymous):

Okay, first of all, you are multiplying three things... \[\large s\]\[\large s\]and\[\frac12s\] So, what's their product? And for heaven's sake, s^2 times 0.5 is not s, you just treated s^2 as 2s, lol

OpenStudy (anonymous):

Cat got your tongue, Mr. Man? What IS your name, anyway? ^.^ And if you want, I'll give a hint :)

OpenStudy (anonymous):

My name is Austin. I'm racking my brain, but I'll take a hint if you would. If you could believe it i'm actually pretty good at math.

OpenStudy (anonymous):

You're pretty good in maths? Of course I believe that :D You're like, way older than I am, after all XD And we all have our off-days Now, when you multiply two variables, just add their exponents. Like this: \[\Large n \times n = n^2\] When you multiply them with numbers, just attach the number... \[\Large 5\times m = 5m\] Combine them, \[\Large 5p \times 4p = (5\times4)p^2=20p^2\] That being said, what's the product of... \[\Large s \times s \times \frac12s = 1s\times 1s\times \frac12s=\color{red}?\]

OpenStudy (anonymous):

oh, and when a variable has no exponent, it's understood that the exponent is 1. that should give you a good start at things :)

OpenStudy (anonymous):

You have a nasty habit of disappearing and or staying eerily quiet... You're not a wall, are you? ^.^

OpenStudy (anonymous):

Nice hint(: No wall here haha. I just feel dumb i know how simple that little equation is, but it just isnt clicking.

OpenStudy (anonymous):

okay, one at a time... you multiply 3 s's right? so you should have s^3 there somewhere. (obviously. all your answer choices involve s^3, did you notice? XD) right?

OpenStudy (anonymous):

YEa i got that.

OpenStudy (anonymous):

Okay, so you have s^3 and you multiply their coefficients...1, 1, and 1/2 What's the product?

OpenStudy (anonymous):

.5

OpenStudy (anonymous):

Stick to 1/2 It's the same. So the product of s, s, and 1/2s was simply \[\Large \frac12s^3\] after all.

OpenStudy (anonymous):

(: 1/2 s^3 seemed to simple.

OpenStudy (anonymous):

yup. so what's your final answer?

OpenStudy (anonymous):

1/2 s^3 ?

OpenStudy (anonymous):

You wish. That's not even in your choices XD

OpenStudy (anonymous):

You weren't listening to me... were you?

OpenStudy (anonymous):

That's what i was thinking. I'm still rather lost. What you say that i missed?

OpenStudy (anonymous):

did*

OpenStudy (anonymous):

hmm... look back. the very first thing I told you. read it again :D

OpenStudy (anonymous):

The volume of a pyramid is equal to one third of the product of its length, width, and height. So i would now multiply (1/3)(1/2 s^3).

OpenStudy (anonymous):

Righty-oh :3

OpenStudy (anonymous):

Today just isn't my day. Ok the answer is (B).

OpenStudy (anonymous):

That wasn't so bad, was it? ^.^

OpenStudy (anonymous):

Woooo you can say that again!! Thank you so much @PeterPan

OpenStudy (anonymous):

What's with the tagging? I'm already here, silly. And while I'm here, just call me Peter, *Austin* (that or Pan... that's what my friends call me ^.^) And for goodness' sake, practice XD

OpenStudy (anonymous):

Definately Pan (: I'll be practicing can't get showed up like that again.

OpenStudy (anonymous):

I'll "motivate" you, shall I? I'm 14 XD

OpenStudy (anonymous):

And truth be told, this looks like a problem I'd have been given two years ago... are you getting rusty or something? XD

OpenStudy (anonymous):

I just pushed off some of my math credits for too long. Need to get my priority's straight.(:

OpenStudy (anonymous):

it's 'priorities' :P What have you been doing, I wonder? Is your head floating someplace else? XD

OpenStudy (anonymous):

Hense the name.(: Did you get my testimonial?

OpenStudy (anonymous):

Of course I did. How very touching :) thanks ^.^

OpenStudy (anonymous):

Anytime(: